# Hyperbolic embedding of brain networks as a tool for epileptic seizures forecasting

### v2

@@ I Introduction
Previous studies [12, 13] showed that brain connectivity matrices, estimated from intracranial electroencephalography (EEG) recordings, allowed the identification of preictal states. In the present study we used network embedding methods to predict the risk of seizure occurrence. Our approach aimed at exploring hyperbolic geometry to represent functional brain networks of patients with epilepsy, and to work in a more adapted space representation to unveil patterns that could potentially result in robust biomarkers for seizure forecasting.
@@ II.1 Data acquisition
The EEG dataset used in this study was previously presented in [14]. Briefly, daily 10-minute resting-state intracranial EEG recordings were obtained in 10 patients (mean age 30.7 years) with drug-resistant focal epilepsy from January 2019 to July 2021 in the Epilepsy Unit of the Pitié-Salpêtrière Hospital (Paris, France). The study was conducted in accordance with the Helsinki Declaration and approved by an institutional review board (project C11-16 and C19-55 of the French National Institute of Health and Medical Research sponsor).
@@ II.2 Hyperbolic embedding of brain networks
Networks filtering. The initial stage of network embedding consists of filtering the PLV matrices by applying a threshold to cancel a percentage of the weakest values. Here, we followed Ref. [18], and the synchrony matrices were filtered such that the final networks reached a predetermined mean degree, fixed here to 4. Although a spanning tree filter could be applied to ensure a unique acyclic subgraph that connects all nodes [19], here we preferred the method of Ref. [18] as it emphasizes the intrinsic global and local properties of the network, while preserving its sparsity.
@@ II.3 Hyperbolic score disk calculation
$\mathrm{P}_{i}^{\mathrm{ref}}\left({W_{i}^{k}}\right)$ is the hyperbolic Gaussian model of nodes $i$ in the reference interictal networks, estimated at the position of the $i^{\mathrm{th}}$ node of a network $k$. Once the scores of all the nodes have been calculated, the disk is discretized with a grid, and the errors of the nodes can be interpolated in the whole disk (one for each network). To perform the interpolation, the scores of the $n=8$ nearest nodes to each pixel of the grid are averaged, with the weights being inversely proportional to the distances between the pixel and the nodes’ positions on the disk. This results in a Hyperbolic Score Disk (HDS) for each network (see Fig. 1).
@@ III.1 Discrimination of preictal networks
We first evaluated the ability of the proposed method to discriminate between interictal and preictal networks. Classification performances were assessed by AUC, the accuracy, and the F1-score and are presented in Tab. 1. In general, the connectivity patterns associated with activities of separated frequency bands cannot distinguish between preictal and interictal days. Only networks estimated from low $\gamma$ oscillations (30-49 Hz) performed better than a random classifier. Nevertheless, classification performances considerably increased when the information from different frequency bands was combined for each patient: All patients reached AUCs of $\geqslant 0.78$ (mean AUC $=0.93$, mean F1 score $=0.79$, and mean accuracy $=0.85$). These findings suggest that the combination of the connectivity in different frequency bands conveys information that can distinguish preictal from interictal epochs.
@@ III.2 Forecasting of upcoming seizure(s)
Forecasting results were consistent with those of our previous study, in which we demonstrated that brain connectivity obtained from vigilance-controlled resting-state recordings could accurately predict the risk of upcoming seizures. [14]. The results, presented in Table. 2, demonstrate the forecasting performance obtained for the distinct frequency bands when the prediction models for each day were trained with data from all preceding days. Combining information from different frequency bands yielded a good predictor with an average Brier score and BSS of $0.12$ [min. = 0.003, max. = 0.48] and $0.36$ [min. = -0.43, max. = 0.7O], respectively. These values confirm that this prediction model performed better than a non-informative model to forecast the risk of upcoming seizure(s). We notice, however, that connectivity networks from the $\beta$ and low $\gamma$ band, also provided good predictors, in agreement with previous findings [14].
@@ IV Discussion
The classification performances demonstrated a high degree of discriminating power when the information from brain connectivity in different frequency bands was combined for each patient. The combination of networks from different frequency bands yielded the most accurate predictions (an accuracy of 85% and F1-score of 79%), but it is notable that connectivity networks from the beta and low gamma bands also demonstrated excellent forecasting results. For all patients, the predictive performance was enhanced by increasing the quantity of learning data. For each day, the optimal prediction was achieved when the models were recursively trained with networks from all preceding days. For these models, false positives were only detected in two patients, and only occurred immediately after consecutive preictal periods. This may indicate a pro-ictal condition, in which seizures are more likely to occur after a day of seizures. However, homeostatic control mechanisms could prevent their emergence.
It should be noted that other non-Euclidean embeddings (e.g. spherical or elliptic), and that alternative dimensional reduction methods (e.g., Isomaps, Locally Linear embeddings [28]) can also be used. Similarly, representations in different hyperbolic spaces, such as Lorentz-Klein, $d-$dimensional Poincaré’s ball, or the half-space model, can be used as alternatives to Poincaré’s disk. Although other hyperbolic mappings (e.g., Mercator [25] HyperMap [34] or Hydra [35] among others) are alternatives to projecting brain networks into the 2D Poincaré disk, the coalescent embedding method used here encompasses various benefits. These include a short computational time and the absence of stochastic minimization procedures, thus preventing errors introduced by local minima.

### v3

@@ I Introduction
Existing methods for predicting epileptic seizures primarily depend on long-term electroencephalographic (EEG) recordings, whether from the scalp or intracranial sources. This approach places considerable demands on both data collection and processing [12, 13]. These models are designed to identify specific preictal changes associated with each seizure, without accounting for alterations caused by shifts in vigilance states. However, such methods necessitate either real-time detection of vigilance states or the integration of various reference interictal states. While some methods achieve high sensitivity, they are often associated with high false positive rates, resulting in frequent false alarms that can diminish their clinical utility [14, 15]. Additionally, performance varies significantly across studies due to differences in datasets, feature selection, and model architectures, underscoring the absence of a standardized, generalizable solution for seizure prediction. Despite these challenges, recent advancements in brain network analysis have created new opportunities for developing more efficient and interpretable predictive models.
Previous studies [16, 17] have shown that brain connectivity matrices derived from intracranial EEG data can effectively identify preictal states. Building on these findings, we introduce a novel seizure forecasting approach that addresses key limitations of existing methods. Unlike continuous monitoring systems, our approach focuses on two main objectives: (1) determining whether patient-specific EEG connectivity during vigilance-controlled resting-state periods differs between seizure and non-seizure days, and (2) prospectively evaluating its relevance for daily seizure risk forecasts using calibrated predictive models. To estimate seizure risk, we utilize a hyperbolic network embedding method to represent patients’ functional connectivity in a low-dimensional space. This representation leverages hierarchical, small-world, and other key properties of brain networks, offering a more suitable framework for capturing connectivity patterns associated with seizure occurrence. Our method outperforms various forecasting models, including traditional machine learning approaches. By identifying robust biomarkers for seizure forecasting, our approach requires less training data while delivering superior performance, paving the way for practical and efficient solutions in epilepsy management.
@@ II.1 Data acquisition
The EEG dataset used in this study was previously presented in [18]. Briefly, daily 10-minute resting-state intracranial EEG recordings were obtained in 10 patients (mean age 30.7 years) with drug-resistant focal epilepsy from January 2019 to July 2021 in the Epilepsy Unit of the Pitié-Salpêtrière Hospital (Paris, France). The study was conducted in accordance with the Helsinki Declaration and approved by an institutional review board (project C11-16 and C19-55 of the French National Institute of Health and Medical Research sponsor). The implanted brain regions and the number of electrodes vary between patients, with a median of 20 electrodes per patient and a range from 10 to 62. The original study associated with this cohort provides detailed demographic and clinical information for each patient [18].
@@ II.2 Hyperbolic embedding of brain networks
Networks filtering. The initial stage of network embedding consists of filtering the PLV matrices by applying a threshold to cancel a percentage of the weakest values. Here, we followed Ref. [22], and the synchrony matrices were filtered such that the final networks reached a predetermined mean degree, set to 4, as recommended for small-size networks. Although a spanning tree filter could be applied to ensure a unique acyclic subgraph that connects all nodes [23], here we preferred the method of Ref. [22] as it emphasizes the intrinsic global and local properties of the network, while preserving its sparsity.
@@ II.3 Hyperbolic score disk calculation
$\mathrm{P}_{i}^{\mathrm{ref}}\left({W_{i}^{k}}\right)$ is the hyperbolic Gaussian model of nodes $i$ in the reference interictal networks, estimated at the position of the $i^{\mathrm{th}}$ node of a network $k$. Once the scores of all the nodes have been calculated, the disk is discretized with a grid, and the errors of the nodes can be interpolated in the whole disk (one for each network). To perform the interpolation, the scores of the $n=8$ nearest nodes to each pixel of the grid are averaged, with the weights being inversely proportional to the distances between the pixel and the nodes’ positions on the disk. This results in a Hyperbolic Score Disk (HSD) for each network (see Fig. 1).
@@ III.1 Discrimination of preictal networks
We first evaluated the ability of the proposed method to discriminate between interictal and preictal networks. Classification performances were assessed by AUC, the accuracy, and the F1-score and are presented in Tab. 1. In general, the connectivity patterns associated with activities of separated frequency bands cannot distinguish between preictal and interictal days. Only networks estimated from low $\gamma$ oscillations (30-49 Hz) performed better than a random classifier. Nevertheless, classification performances considerably increased when the information from different frequency bands was combined for each patient: All patients reached AUCs of $\geqslant 0.78$ (mean AUC $=0.93$, mean F1 score $=0.79$, and mean accuracy $=0.85$). These findings suggest that combining connectivity across different frequency bands provides valuable information for distinguishing preictal from interictal epochs. Notably, the hyperbolic embedding method outperformed a standard machine learning model, specifically a support vector machine (SVM). The SVM model achieved an average AUC of $0.79$, an F1 score of $0.78$, and a mean accuracy of $0.77$, highlighting the superior performance of the hyperbolic embedding approach.
@@ III.2 Forecasting of upcoming seizure(s)
Forecasting results were consistent with those of our previous study, in which we demonstrated that brain connectivity obtained from vigilance-controlled resting-state recordings could accurately predict the risk of upcoming seizures. [18]. The results, presented in Table. 2, demonstrate the forecasting performance obtained for the distinct frequency bands when the prediction models for each day were trained with data from all preceding days. Combining information from different frequency bands yielded a good predictor with an average Brier score and BSS of $0.12$ [min. = 0.003, max. = 0.48] and $0.36$ [min. $=-0.43$, max. $=0.70$], respectively. In comparison, the SVM model achieved a mean Brier score of $0.13$. It is worthy of notice that connectivity networks from the $\beta$ and low $\gamma$ band also provided good predictors, in agreement with previous findings [18]. These results demonstrate that our prediction model outperformed other approaches, including a standard machine learning model, in forecasting the risk of upcoming seizure(s).
@@ IV Discussion
The classification performances demonstrated a high degree of discriminating power when the information from brain connectivity in different frequency bands was combined for each patient. The combination of networks from different frequency bands yielded the most accurate predictions (an accuracy of 87% and F1-score of 89%), but it is notable that connectivity networks from the beta and low gamma bands also demonstrated excellent forecasting results. For all patients, the predictive performance was enhanced by increasing the quantity of learning data. For each day, the optimal prediction was achieved when the models were recursively trained with networks from all preceding days. For these models, false positives were only detected in two patients, and only occurred immediately after consecutive preictal periods. This may indicate a pro-ictal condition, in which seizures are more likely to occur after a day of seizures. However, homeostatic control mechanisms could prevent their emergence.
Unlike current models that rely on long-term EEG recordings to track changes preceding seizures, our findings indicate that short resting-state recordings can offer valuable insights into the preictal state on a daily scale. It is worth noting that a direct comparison to traditional seizure prediction methods is not entirely feasible due to methodological differences, particularly in the duration of recordings utilized. However, compared to alternative models such as the SVM classifier [18] or a simple anteriority-based model, our approach exhibits superior performance. Additionally, most EEG-based methods for seizure prediction provide forecasts within minutes to, at most, an hour before a seizure event [14, 15]. In contrast, our method provides a single daily forecast, offering an extended time window for intervention or monitoring throughout the day.
It should be noted that other non-Euclidean embeddings (e.g. spherical or elliptic), and that alternative dimensional reduction methods (e.g., Isomaps, Locally Linear embeddings [32]) can also be used. Similarly, representations in different hyperbolic spaces, such as Lorentz-Klein, $d-$dimensional Poincaré’s ball, or the half-space model, can be used as alternatives to Poincaré’s disk. Although other hyperbolic mappings (e.g., Mercator [29] HyperMap [38] or Hydra [39] among others) are alternatives to projecting brain networks into the 2D Poincaré disk, the coalescent embedding method used here encompasses various benefits. These include a short computational time and the absence of stochastic minimization procedures, thus preventing errors introduced by local minima. Similarly, while alternative methods based on the Minimum Spanning Tree (MST) can be used for filtering connectivity matrices [40], the MST [23] and MST+ECO [22] approaches exhibit poor performance in forecasting preictal states, yielding an average accuracy $\leq 0.56$ and a F1-score $\leq 0.4$. These results highlight the advantages of our chosen filtering method.